OpenStudy (anonymous):

Which statements are true? ***CHoose ALL that apply!! A. An angle that measures pi/3 radians also measures 45 degrees. B. An angle that measures pi/6 radians also measures 30 degrees. C. An angle that measures 180 degrees also measures pi radians. D. An angle that measures 30 degrees also measures pi/3 radians. please explain?? and its choose ALL that apply! :)

OpenStudy (anonymous):

hey @zepdrix :) thanks for coming! do u get this?

OpenStudy (anonymous):

pi = 180 degree. A) 180 / 3 isn't = 45 wrong B) 180 / 6 = 30 correct C) 180 = pi correct D) 180 / 3 isn't = 60 wrong

OpenStudy (anonymous):

ohh okay so that's all i had to do? divide them with 180 to see if they matched?

OpenStudy (anonymous):

and 180 always represents pi yes? @nphuongsun93 :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i see.. thanks so much! =)